# MonoPatchNeRF: Improving Neural Radiance Fields with Patch-based Monocular Guidance

### v1

@@ 5 Conclusion
We propose MonoPatchNeRF, a patch-based regularized NeRF model that aims to produce geometrically accurate models. We demonstrate the effective use of monocular geometry estimates with patch-based ray sampling optimization and density constraints, as well as the effectiveness of NCC and SSIM photometric consistency losses between patches from virtual and training views. Our method significantly improves geometric accuracy, ranking top in terms of $F_{1}$, SSIM, and LPIPS compared to state-of-the-art regularized NeRF methods on the challenging ETH3D MVS benchmark.
Acknowledgement This work is supported in part by NSF IIS grants 2312102 and 2020227. SW is supported by NSF 2331878 and 2340254, and research grants from Intel, Amazon, and IBM. Thanks to Zhi-Hao Lin for sharing the code for pose interpolation and video generation.
@@ 6 Training Details
Throughout our experiments, we use $\lambda_{huber},\lambda_{depth},\lambda_{\nabla depth},\lambda_{normal},\lambda_{\nabla normal}$,
$\lambda_{SSIM},\lambda_{NCC},\lambda_{MVS}$ set to $1.0$, $0.05$, $0.025$, $0.001$, $0.0005$, $41E-$ 4, $41E-$ 4, $0.1$. We use L2 norm for $L_{depth}$ and L1 norm for $L_{normal}$. We compute the novel view occludion mask by comparing the angle difference between the rays from the training origin to the training unprojected pixels and the novel view unprojected pixels. Pixels with angles greater than 10 degrees are considered occluded, and are excluded when calculating the novel patches loss $L_{novel}$. We also provide the detailed network architecture in Table 5. Our model contains a total of 16,385,392 parameters and is trained for 50,000 steps on ETH3D in 2.25 hours and 200,000 steps on TanksandTemples in 9 hours using a single Nvidia A40 GPU. We compare our statistics with other baseline models in Table 6.
Table 6: Training and Inference Time. We report the number of steps and time taken for optimization and rendering time for one image with resolution of $640,960$ in ETH3D [29] training scenes. All timings are measured with the same device using single NVidia A40 GPU. The fast training and the rendering time comes from the architectural choice [14].
Model # Steps Training (hh:mm) Inference (ss)
FreeNeRF [34] 200,000 14:26 36.4
RegNeRF [24] 200,000 07:00 30.4
MonoSDF [38] 200,000 20:54 136.0
Neuralangelo [16] 200,000 19:58 61.9
Ours 50,000 02:15 4.6
@@ 7 Baseline Training Details
MonoSDF: We show how varying the initialization bias [atzmon2020sal] of MonoSDF [38] affects its reconstruction quality. We used the author provided configs of TnT on ETH3D, but found that MonoSDF suffer from local minimum in some challenging scenes with original bias parameters due to the scene scale. We therefore contacted the MonoSDF authors and were advised to use a small bias for initialization. Figure 8 compares the reconstruction of MonoSDF given different bias parameters in a challenging scene relief_2 of ETH3D [29].
[…]
Refer to caption Refer to caption
Batch size = 4 Batch size = 16
@@ 9 Additional Qualitative Results
Images: Additional visualization of mesh and novel view synthesis is shown in Figure 11. We provide additional visualizations of our method on subsets of TanksAndTemples [11] advanced scenes in Figure 12 and Figure 13, and on ETH3D [29] in Figure 14 and Figure 15.
Figure 11: Additioanl comparison of novel view images and meshes on ETH3D. We provide additional comparisons with baselines [24, 34, 38, 16].
All our results, except for ones annotated with (Ours-MVS-Depth) in Figure 7 and Table 1 of the main paper, are from our model with monocular cues.

### v2

@@ 5 Conclusion
We propose MonoPatchNeRF, a patch-based regularized NeRF model that aims to produce geometrically accurate models. We demonstrate the effective use of monocular geometry estimates with patch-based ray sampling optimization and density constraints, as well as the effectiveness of NCC and SSIM photometric consistency losses between patches from virtual and training views. Our method significantly improves geometric accuracy, ranking top in terms of $F_{1}$, SSIM, and LPIPS compared to state-of-the-art regularized NeRF methods on the challenging ETH3D MVS benchmark. Still, there are many potential directions for improvement, including: guided sampling of virtual view patches; joint inference of geometry with single-view predictions; including per-image terms to better handle lighting effects; expanding material models, e.g. with both diffuse and specular terms per point; incorporating semantic segmentation; and reducing memory and computation requirements.
Acknowledgement This work is supported in part by NSF IIS grants 2312102 and 2020227. S.W. is supported by NSF 2331878 and 2340254, and research grants from Intel, Amazon, and IBM. Thanks to Zhi-Hao Lin for sharing the code for pose interpolation and video generation, and Bowei Chen and Liwen Wu for proofreading the paper.
@@ 6 Training Details
We provide the detailed network architecture in Table 5. Our model contains a total of 16,385,392 parameters and is trained for 50,000 steps on ETH3D in 2.25 hours and 200,000 steps on TanksandTemples in 9 hours using a single Nvidia A40 GPU.
@@ 7 Baseline Training Details
MonoSDF: We show how varying the initialization bias of MonoSDF [55] affects its reconstruction quality. We used the author provided configs of TnT on ETH3D, but found that MonoSDF suffer from local minimum in some challenging scenes with original bias parameters due to the scene scale. We therefore contacted the MonoSDF authors and were advised to use a small bias for initialization. Figure 8 compares the reconstruction of MonoSDF given different bias parameters in a challenging scene relief_2 of ETH3D [36].
Figure 8: Visualization of MonoSDF with different paramters in Relief_2. MonoSDF1.0 denotes MonoSDF trained with the default bias parameter (1.0) in the code provided by the authors on the large scale TnT [15] evaluation. MonoSDF0.1 denotes MonoSDF trained with the parameter suggested by the authors (0.1) for large-scale scenes specifically. MonoSDF [55] reconstructs better mesh given smaller bias, and falls into local minimum with original bias.
Neuralangelo: For Neuralangelo [20] experiment on ETH3D [36], we follow author provided setup on TanksAndTemples [15], but use a batch size of 4 instead of 16 to run on the same device settings. We additionally disable image embedding features, as we empirically found it to worsen the results. One visualization of results with different batch size is present in Figure 9. The $F$-score2cm, $F$-score5cm of high and low batch size results are 1.53, 11.5 and 1.46, 11.8.
@@
8 Depth Evaluation
We provide evaluation of depth maps from radiance based methods and MVS methods in Tab. 6 following robust multi-view benchmark [37]. We report the Absolute Relative Error (rel) and Inlier Ratio ($\tau$) with a threshold of 1.03, and split the evaluated methods based on the requirement for poses, depth ranges, and intrinsics. To better compare the performance, we additionally group NeRF models together. We show that though not as accurate as MVS methods, we predict more complete depths, as our rel is lower than all classical MVS methods.
Table 6: Multi-view depth evaluation with different settings: a) Classical approaches; b) with poses and depth range, without alignment; c) absolute scale evaluation with poses, without depth range and alignment; d) without poses and depth range, but with alignment; e) neural radiance field based models. ‘med‘ means alignment based on median ground truth depth and the median predicted depth. Numbers with are from our results, while others are from DUSt3R [44]. The best results for each setting are in bold.
Methods GT GT GT Align ETH3D
Pose Range Intrinsics rel $\downarrow$ $\tau\uparrow$
(a) COLMAP [33, 35] $\checkmark$ $\times$ $\checkmark$ $\times$ 16.4 55.1
COLMAP Dense [33, 35] $\checkmark$ $\times$ $\checkmark$ $\times$ 89.8 23.2
ACMMP [47] $\checkmark$ $\times$ $\checkmark$ $\times$ 16.0 91.6
(b) MVSNet [51] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 35.4 31.4
MVSNet Inv. Depth [51] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 21.6 35.6
Vis-MVSSNet [56] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 10.8 43.3
MVS2D ScanNet [49] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 27.4 4.8
MVS2D DTU [49] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 99.0 11.6
(c) DeMon [40] $\checkmark$ $\times$ $\checkmark$ $\times$ 19.0 16.2
DeepV2D KITTI [38] $\checkmark$ $\times$ $\checkmark$ $\times$ 30.1 9.4
DeepV2D ScanNet [38] $\checkmark$ $\times$ $\checkmark$ $\times$ 18.7 28.7
MVSNet [51] $\checkmark$ $\times$ $\checkmark$ $\times$ 507.7 8.3
MVSNet Inv. Depth [51] $\checkmark$ $\times$ $\checkmark$ $\times$ 60.3 5.8
Vis-MVSNet [56] $\checkmark$ $\times$ $\checkmark$ $\times$ 51.5 17.4
MVS2D ScanNet [49] $\checkmark$ $\times$ $\checkmark$ $\times$ 30.7 14.4
MVS2D DTU [49] $\checkmark$ $\times$ $\checkmark$ $\times$ 78.0 0.0
Robust MVD Baseline [37] $\checkmark$ $\times$ $\checkmark$ $\times$ 9.0 42.6
(d) DeMoN [40] $\times$ $\times$ $\checkmark$ $\|\mathbf{t}\|$ 17.4 15.4
DeepV2D KITTI [38] $\times$ $\times$ $\checkmark$ med 27.1 10.1
DeepV2D ScanNet [38] $\times$ $\times$ $\checkmark$ med 11.8 29.3
DUSt3R 224-NoCroCo[44] $\times$ $\times$ $\times$ med 9.51 40.07
DUSt3R 224[44] $\times$ $\times$ $\times$ med 4.71 61.74
DUSt3R 512[44] $\times$ $\times$ $\times$ med 2.91 76.91
(e) RegNeRF [28] $\checkmark$ $\times$ $\checkmark$ $\times$ 24.9 15.0
FreeNeRF [48] $\checkmark$ $\times$ $\checkmark$ $\times$ 194.0 7.3
Ours $\checkmark$ $\times$ $\checkmark$ $\times$ 7.4 67.2
Ours (MVS-Depth) $\checkmark$ $\times$ $\checkmark$ $\times$ 7.2 83.5
9 Sparse-view Tanks and Temples Comparison
We compare our method with Neuralangelo [20] on Tanks and Temples [15] with 1/5 views that are uniformly sampled from original views in Tab. 7. We follow Neuralangelo to preprocess the scene, and use the same parameters (include the batch size) as the original paper except for using 200,000 instead of 500,000 training steps due to reduced input images. Our method outperform Neuralangelo in all scenes, showing that our method works better for the challenging sparse view setup. See Fig 10 for qualitative comparison.
Table 7: Comparison on TnT [15] with sparse input views. We report the $F$-score of our method and Neuralangelo [20] in three large-scale TnT scenes following Neuralangelo preprocessing. Best results are in bold.
Meetingroom Courthouse Barn
Ours 14.0 16.0 30.8
Neuralangelo 1.7 7.7 5.7
Figure 10: Training view visualization on sparse view TnT. We compare the rendered color images, depth map, and normals map from the training view on Meetingroom for our method and Neuralangelo [20].
@@ 11 Additional Qualitative Results
Images: We present one additional compairson like our teaser figure in Figure 12. Additional visualization of mesh and novel view synthesis are shown in Figure 13. We also provide additional visualizations of our method on subsets of TanksAndTemples [15] advanced scenes in Figure 14 and Figure 15, and on ETH3D [36] in Figure 16 and Figure 17.
Videos: We present a free-view rendering of scenes Relief_2, Facade, and Kicker on ETH3D [36] with trajectory interpolated from training poses. We also provide comparisons with RegNeRF [28], Neuralangelo[20], and MonoSDF [55]. There are only 31, 76, 31 training views for the scenes, but the rendering is realistic and the video is smooth. All our results, except for ones annotated with (Ours-MVS-Depth) in Figure 7 and Table 1 of the main paper, are from our model with monocular cues.
Figure 12: Additional qualitative comparison on kicker. We provide additional test view comparisons with baselines [28, 48, 55, 20].
Figure 13: Additional comparison of novel view images and meshes on ETH3D. We provide additional comparisons with baselines [28, 48, 55, 20].